I cannot figure out the formula...If I invest $21,134, part @ 8% and 6%. The amount made @ 8% is $478.88 more than the amount @ 6%. What are the amounts invested @ 8% and 6%
Answer by mananth(16946)

21134 @ 8% & 6 %
let x be the investment on 6%
21134-x will be @8%
..
0.08(21134-x)=0.06x + 478.88
1690.72-0.08x=0.06x+478.88
1690.72-478.88=0.06x+0.08x
1211.84 = 0.14x
x= $ 8656 @6%
21134-8656= 12,478 @ 8 %